OpenStudy (anonymous):

I need to graph a linear equation, -2x + 3y = 0, I know that for A) the points are 0, 0 but the B points I am having a time with. Need help!!!!

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

-2x+3y=0 3y=2x y=2x/3 so you can put in any x value and find the corresponding y value.

OpenStudy (anonymous):

For example: if x=1 then y=2(1)/3 y=2/3 if x=2 then y=2(2)/3 y=4/3

OpenStudy (anonymous):

when you connect the points you should have a straight line. dont forget to use graph paper.

OpenStudy (anonymous):

On the 0,0 point lets check: y=2x/3 if x=0 then y=2(0)/3 anything multiplied by 0 is 0. so y=0/3 0 divided by 3 is 0. so you are correct.